CREAM Flight 2008
CREAM-IV at McMurdo, Nov. 15th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 15th of November at McMurdo.
- Before starting the CPT, Prashant tested the Chercam reset code by running for 2 hours in flight mode and confirmed that the Chercam reset is working so no more of continuous event number mismatching errors in Chercam.
- Also updated script files are uploaded into the SFC1 before the SFC1 CPT started.
CREAM-IV at McMurdo, Nov. 14th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 14th of November at McMurdo.
- For most of day, TCD/CD data were taken with different threshold values for the calibration of the threshold values.
CREAM-IV at McMurdo, Nov. 13th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 13th of November at McMurdo.
- The TCD dead time debugging continued with threshold value increase in the ZHi trigger. Then the even number mismatch error reduced dramatically and the live time increased to about 90% in flight mode. To understand the correct ZHi threshold values for the flight, Tyler has asked Nick/Isaac. This is done with the default cdaq code.
CREAM-IV at McMurdo, Nov. 12th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 12th of November at McMurdo.
- TCD problem with the PMT#4/5 HV resolved by reconnecting the 12 V power line to the HV PS which was cut during the HSK circuit modification.
- After that the hkconv4.dat file modified to have correct value in TCDPwr12. The new hkconv4.dat file will be updated into the relevant computers.
CREAM-IV at McMurdo, Nov. 11th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 11th of November at McMurdo.
- The CREAM instrument is now on its legs after removing the moving table that supported it.
- Solar panels preparation work started in the afternoon when the wind becomes lighter. But still it was a rather windy day.



